# Towards Interpretability Without Sacrifice: Faithful Dense Layer Decomposition with Mixture of Decoders

### v1

@@ B.5 Sparse probing
Here, we follow the SAEBench [19] evaluation protocol. In this ‘sample-level’ setting, each text string is labeled with a binary concept at a global level (e.g., the language of the snippet, or its sentiment). This is in contrast to what we refer to as ‘token-level probing’, where each token within the text samples is labeled individually (e.g., whether a word is a certain part of speech). We perform experiments on a total of $24$ sample-level sparse probing tasks with the same ‘maximum mean difference’ feature filtering applied in [19]. The details of the datasets used are summarized in Table 4.
@@
We also explore sparse probing for $10$ features defined at the token-level. For this, we follow [49], and include experiments training probes on the mean feature activations under tokens spanning the surnames of the individuals. We note that this is a significantly harder task, and makes even stronger assumptions about the features the dataset includes, but is nonetheless some additional weak evidence about the relative feature-learning abilities of the sparse models. Through various surnames, we probe for $6$ occupations of individuals, whether or not individuals are alive, and individuals’ labeled gender. We also experimented with probing for compound words as in [49], but found no predictive features in our trained models. Details of the surname token-level probing datasets (and the total training examples the tokenizers could parse) are included in Table 5.
@@
Binary probes are trained on $80\%$ of the training data (randomly shuffled) with the sklearn library’s LogisiticRegression module with parameters:
@@
B.5.1 Sparse probing results
We show in Figure 13 results on $20$ additional (sample-level) sparse probing tasks, where MxDs remain competitive with the baselines. We also plot the expert activation (of the single expert with the highest F1 test set score) for the positive/negative classes for all tasks split across Figures 14 and 15. One can observe a certain degree of separability between the two semantic clusters of data given by the expert coefficient, thus confirming that individual experts are learning to specialize to particular high-level features.
We also include results on $10$ token-level probing tasks in Figure 16, with the corresponding activation densities displayed in Figure 17. Whilst MxDs appear to perform slightly less well here on average, they remain competitive as expected.
@@ B.6 Ablations
B.6.1 Choice of sparsity constraint
We first train a variety of MxDs on GPT2 models with the TopK activation function [23] and instead train models with a ReLU followed by an explicit $\lambda||.||_{1}$ sparsity penalty on the specialized components in addition to the reconstruction loss [22]. We show the results in Figure 18, where, similarly to [26], we find the TopK activation to dominate on the sparsity-accuracy frontier–we thus use the TopK activation for all experiments.
@@ B.6.1 Choice of sparsity constraint
B.6.2 Choice of MxD encoder
Secondly, we show in Figure 19 the benefits of MxDs’ flexibility in inheriting the original MLP layer’s encoder form/activation function. All models here are trained from scratch for the same number of tokens and with the same experimental setup as in Section 3.1, with $K=32$. In the first 3 left-most subfigures, we see the Normalized MSE is as low as half when using GELU vs the non-native ReLU activation.
We next ablate the impact of inheriting the same encoder as the Llama-3.2-3B base model. In the rightmost subfigure of Figure 19, we train MxDs with ReLU-MLP, GELU-MLP, and Swish-GLU encoders. As can be seen, using a GLU with a Swish activation model (matching the base model architecture) yields a Normalized MSE almost an order of magnitude smaller than MLPs with GELU/ReLU.
@@ C.1 Expert/feature balance
Following the code of [27, 99], we log how often each unit of specialism/feature is used, over a fixed window of $\sim 1$M tokens. We show in Figure 20 the feature frequency at the end of training, where we observe that MxDs see a similar healthy balance of experts to the frequency of usage of features in the baselines.
@@ C.2 Shared experts
We train two sets of models with a base $K=32$ on GPT2-small and pythia-410m, using $a:=2$. We first show in Figure 21 the indices of the top-activating experts for the 2 model variants on a template prompt, after training has finished. On the left-hand side of Figure 21, the models route all tokens through the same shared expert at position 1. However, we see on the right-hand side that training with the ‘random-K’ strategy breaks the dependence on a shared expert in position 1. Furthermore, we include in Figure 22 the corresponding train-time MSE loss for the 4 models here as ablations–observing that the random-K strategy also brings comparable performance. Based on these experiments, we recommend this simple training strategy if one desires MxD models without shared experts.
@@ Appendix D Detailed experimental setup
[…]
@@ Appendix E Additional qualitative results
We show in Figures 24 and 23 tokens activating the first 9 experts as they appear numerically. We sample 6 bins of expert coefficient value to show both tokens that highly activate the experts and those that do so only mildly. As can be seen, both high- and low-level specializations emerge in both GPT and Pythia models.
@@
Figure 23: Tokens activating the first $9$ numerical experts on MxDs with $K=32$ trained on Pythia-410m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to MMO games, abbreviations, and words in specific contexts.
Figure 24: Tokens activating the first $9$ numerical experts on MxDs with $K=32$ trained on GPT2-124m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to punctuation, names, and months.

### v2

@@ B.7 Sparse probing
Here, we follow the SAEBench [19] evaluation protocol. In this ‘sample-level’ setting, each text string is labeled with a binary concept at a global level (e.g., the language of the snippet, or its sentiment). This is in contrast to what we refer to as ‘token-level probing’, where each token within the text samples is labeled individually (e.g., whether a word is a certain part of speech). We perform experiments on a total of $24$ sample-level sparse probing tasks with the same ‘maximum mean difference’ feature filtering applied in [19]. The details of the datasets used are summarized in Table˜7.
@@
We also explore sparse probing for $10$ features defined at the token-level. For this, we follow [49], and include experiments training probes on the mean feature activations under tokens spanning the surnames of the individuals. We note that this is a significantly harder task, and makes even stronger assumptions about the features the dataset includes, but is nonetheless some additional weak evidence about the relative feature-learning abilities of the sparse models. Through various surnames, we probe for $6$ occupations of individuals, whether or not individuals are alive, and individuals’ labeled gender. We also experimented with probing for compound words as in [49], but found no predictive features in our trained models. Details of the surname token-level probing datasets (and the total training examples the tokenizers could parse) are included in Table˜8.
@@
Binary probes are trained on $80\%$ of the training data (randomly shuffled) with the sklearn library’s LogisticRegression module with parameters:
@@
B.7.1 Sparse probing results
We show in Figure˜12 results on $20$ additional (sample-level) sparse probing tasks, where MxDs remain competitive with the baselines. We also plot the expert activation (of the single expert with the highest F1 test set score) for the positive/negative classes for all tasks split across Figures˜13 and 14. One can observe a certain degree of separability between the two semantic clusters of data given by the expert coefficient, thus confirming that individual experts are learning to specialize to particular high-level features.
We also include results on $10$ token-level probing tasks in Figure˜15, with the corresponding activation densities displayed in Figure˜16. Whilst MxDs appear to perform slightly less well here on average, they remain competitive as expected.
@@ B.8 Ablations
B.8.1 Choice of sparsity constraint
We first train a variety of MxDs on GPT2 models with the TopK activation function [23] and instead train models with a ReLU followed by an explicit $\lambda||.||_{1}$ sparsity penalty on the specialized components in addition to the reconstruction loss [22]. We show the results in Figure˜17, where, similarly to [26], we find the TopK activation to dominate on the sparsity-accuracy frontier–we thus use the TopK activation for all experiments.
@@ B.8.1 Choice of sparsity constraint
B.8.2 Choice of MxD encoder
Secondly, we show in Figure˜18 the benefits of MxDs’ flexibility in inheriting the original MLP layer’s encoder form/activation function. All models here are trained from scratch for the same number of tokens and with the same experimental setup as in Section˜3.1, with $K=32$. In the first 3 left-most subfigures, we see the Normalized MSE is as low as half when using GELU vs the non-native ReLU activation.
We next ablate the impact of inheriting the same encoder as the Llama-3.2-3B base model. In the rightmost subfigure of Figure˜18, we train MxDs with ReLU-MLP, GELU-MLP, and Swish-GLU encoders. As can be seen, using a GLU with a Swish activation model (matching the base model architecture) yields a Normalized MSE almost an order of magnitude smaller than MLPs with GELU/ReLU.
@@ C.1 Expert/feature balance
Following the code of [27, 107], we log how often each unit of specialism/feature is used, over a fixed window of $\sim 1$M tokens. We show in Figure˜19 the feature frequency at the end of training, where we observe that MxDs see a similar healthy balance of experts to the frequency of usage of features in the baselines.
@@ C.2 Shared experts
We train two sets of models with a base $K=32$ on GPT2-small and pythia-410m, using $a:=2$. We first show in Figure˜20 the indices of the top-activating experts for the 2 model variants on a template prompt, after training has finished. On the left-hand side of Figure˜20, the models route all tokens through the same shared expert at position 1. However, we see on the right-hand side that training with the ‘random-K’ strategy breaks the dependence on a shared expert in position 1. Furthermore, we include in Figure˜21 the corresponding train-time MSE loss for the 4 models here as ablations–observing that the random-K strategy also brings comparable performance. Based on these experiments, we recommend this simple training strategy if one desires MxD models without shared experts.
Expert diversity/collapse
To what extent is the shared expert simply learning the base decoder itself? In an initial attempt to study this, we take the same MxD trained with $K=64$ on GPT2, and tabulate the Frobenius norm of differences between the original model’s decoder $\mathbf{D}^{*}$, our learned base decoder $\mathbf{D}$, and the “shared expert” weight matrix $\mathbf{W}_{\text{shared}}=\mathbf{D}\,\text{diag}(\mathbf{c}_{\text{shared}})$ (with index $\text{shared}=19772$), all of which are matrices of shape $(H\times O)$. As can be seen in Table˜9, their difference is significant, showing that the shared expert does not simply learn the original decoder.
As discussed in the limitations section, we highlight that MxDs as formulated contain no explicit loss terms to discourage expert collapse, or encourage expert balance or diversity. Whilst we find no need for these in our experiments, they may prove useful (or even necessary) in other settings or future work.
Table 9: Distance between learned shared expert, base decoder, and original decoder
$\left\lVert\mathbf{W}_{\text{shared}}-\mathbf{D}^{*}\right\rVert_{F}$ $\left\lVert\mathbf{D}-\mathbf{D}^{*}\right\rVert_{F}$
$\left(\left|\mathbf{W}_{\text{shared}}-\mathbf{D}^{*}\right|\right)_{:3,:3}$
(9 elements of the difference matrix)
208.08 221.6934 $\begin{bmatrix}0.09&0.03&0.06\\ 0.21&0.12&0.07\\ 0.17&0.15&0.11\end{bmatrix}$
@@ Appendix D Detailed experimental setup
We list in Table˜10 the resources used for each experiment: the GPU and the indicative run-time for a single model. The mlp_expansion_factor column refers to the expansion factor applied to the input dimension to generate the MLP width in the sparse layers (i.e. $H:=I\cdot\text{mlp\_expansion\_factor}$).
@@ Appendix E Additional qualitative results
We show in Figures˜23 and 22 tokens activating the first 9 experts as they appear numerically. We sample 6 bins of expert coefficient value to show both tokens that highly activate the experts and those that do so only mildly. As can be seen, both high- and low-level specializations emerge in both GPT and Pythia models.
@@
Figure 22: Tokens activating the first 99 numerical experts on MxDs with $K=32$ trained on Pythia-410m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to MMO games, abbreviations, and words in specific contexts.
Figure 23: Tokens activating the first 99 numerical experts on MxDs with $K=32$ trained on GPT2-124m; we sample 6 bands of activations to show both tokens that highly activate experts and those that activate them only mildly. Magnitude of activation is denoted by the orange highlight. Moderate specialism emerges, e.g., to punctuation, names, and months.